### v2

@@ Abstract
The major contribution of the current work is to provide a prediction augmented algorithmic framework for balancing the consistency and robustness over strategic agents with non-uniform weights. In particular, through a reduction technique that identifies a subset of representative instances and maps the other given locations to the representative ones, we prove that there exists a strategyproof mechanism achieving a bounded consistency guarantee of $\frac{\sqrt{(1+c)^{2}W^{2}_{\min}+(1-c)^{2}W^{2}_{\max}}}{(1+c)W_{\min}}$ and a bounded robustness guarantee of $\frac{\sqrt{(1-c)^{2}W^{2}_{\min}+(1+c)^{2}W^{2}_{\max}}}{(1-c)W_{\min}}$ in weighted settings, where $c$ can be viewed as a parameter to make a trade-off between the consistency and robustness and $W_{\min}$ and $W_{\max}$ denote the minimum and maximum agents’ weight. We also proved that there is no strategyproof deterministic mechanism that reach $1$-consistency and $O\left(n\cdot\frac{W_{\max}}{W_{\min}}\right)$-robustness in weighted FLP, even with fully predictions of all agents.
@@ 1 Introduction
Facility Location Problem (FLP) is a classic problem in operations research, computer science, and economics [8]. Its goal is to choose one or more facility locations within a given space to optimize a global cost function between users (referred to as agents) and facilities. If only one facility needs to place, we say it single FLP [^1] If not mentioned on purpose, in the following text, we say the FLP as single FLP by default.. In the problem, there are $n$ agents in a metric space $\mathcal{M}$, and each agent $i$ has its preferred location $p_{i}$ and must report a location $\hat{p}_{i}\in\mathcal{M}$ of the facility. We are required to collect this information to choose where to locate the facility $f\in\mathcal{M}$. Once we choose the facility location $f$, the individual cost of each agent $i$ is given by the distance $d(p_{i},f)$ between $p_{i}$ and $f$, where $d(\cdot)$ is the distance function defined on $\mathcal{M}$. Our objective is to optimize a global function $g(P;f)$, where $P$ is the location profile of all the agents. Naturally, here comes to a question: What kind of global cost function we should consider? Motivated by various applications in practical scenarios (see [8] for a more detailed overview), researchers in this area are often interested in the global objectives of (1) egalitarian social cost and (2) utilitarian social cost, where egalitarian social cost is to the maximum cost over all the agents, and utilitarian social cost is the average cost. The study of the former is more difficult as applying predictions on any existing strategyproof mechanisms is hard, for the mechanisms focus on the bounding box of $P$ that prediction can hardly get involved. So the current work is devoted to the study of the latter.
@@
By our assumption, agents are selfish, so an agent $i$ may misreport her preferred location to minimize her individual cost, which means the reported location $\hat{p}_{i}$ is not necessarily equivalent to her real preferred location $p_{i}$. This leads to the strategic facility location problem, where the focus shifts from purely optimizing a cost function to designing mechanisms that are strategyproof, i.e., mechanisms in which agents have no incentive to tell lie about their true preferences [24, 6]. This question was first studied by Procaccia and Tennenholtz [24], focusing on the egalitarian social cost, and Dekel [11] studied to optimize the utilitarian social cost. Prior work in this field has yielded fruitful results. Procaccia and Tennenholtz [24] proved that no deterministic and strategyproof mechanism can achieve better than a 2-approximation, when the goal is the egalitarian social cost, and Meir [21] proposed the coordinate median (CM) mechanism that reaches $\sqrt{n}$-approximation in $n$-dimensional case while optimizing the utilitarian social cost. These results established the framework for subsequent research.
The framework shown above has been applied in establishment of public services such as schools and hospitals, and really made sense [16, 10]. Existing work is often established on the unweighted assumption on the agents, which means that the importance of all agents in this issue is completely equal. This assumption works when the goal is global social well-being, but it often does not hold in real life. For example, in the healthcare field, the elderly have a higher incidence rate. If the healthcare facility is too far from them, their mortality rate will increase significantly [2, 22]. So, the elderly have a greater weight than others in this case. The assumption that agents have weights also has practical significance in the fields like logistics planning [5], disaster assistance [7] and so on. Unweighted assumption cannot deal with these problems well. That’s why strategic FLP with weighted values occurs. In weighted settings, we assume that each agent has a weight value that quantifies its importance to the problem: agents with higher weights take greater social cost and vice versa.
@@
In strategic FLP, there is another great problem remaining: the work mostly focuses on the worst case, which occurs hardly in real practice [1, 9]. That’s because decision makers don’t know some of the information, so if we can predict the missing part, we can take advantage of it in some way. Prediction-augmented algorithms import external predictive data into decision making, allowing decision makers themselves to decide how much they trust the predicted results [14, 13, 21, 30]. When the prediction is accurate, we can trust it more and get a better resulting guarantee than using worst-case analysis, i.e., consistency of the mechanism. On the contrast, when the prediction is arbitrarily wrong, we hope that the wrong prediction does not worsen the result too much and obtains a solid resulting guarantee, i.e., robustness of the mechanism. More precisely speaking, in the framework of prediction-augmented mechanism design of strategic FLP, given the agents’ reported locations $P$, the designer is provided a predicted location $\hat{o}$, regarding the real optimal location $o(P)$. The designer can use the information to make a mechanism, whose result is $f(P,\hat{o})$. Given a social cost function $C$, the mechanism is $\alpha$-consistent when the prediction is accurate $(\hat{o}=o(P))$ if
@@
$\max_{P,\hat{o}}{\{\frac{C(f(P,\hat{o}),P)}{C(o(P),P)}\}}\leq\beta.$
@@ 1.1 Our Contributions
We investigate the prediction-augmented mechanism design of strategic FLP in a weighted setting, where the importance of the agents is unequal according to the weighted values. Our goal is to design deterministic, strategyproof mechanisms that incorporate a prediction of the optimal facility location, and achieve a tunable trade-off between consistency and robustness.
@@
There exists a deterministic and strategyproof mechanism in the weighted FLP with predictions that reaches $\frac{\sqrt{(1+c)^{2}W^{2}_{\min}+(1-c)^{2}W^{2}_{\max}}}{(1+c)W_{\min}}$consistency and $\frac{\sqrt{(1-c)^{2}W^{2}_{\min}+(1+c)^{2}W^{2}_{\max}}}{(1-c)W_{\min}}$robustness.
$W_{\min}$ and $W_{\max}$ denote the minimum and maximum agents’ weight, respectively. The mechanism uses a parameter $c\in[0,1)$ to control the degree of reliance on the prediction. When $c$ is large, the mechanism places more trust in the prediction, achieving better consistency; when $c$ is small, it becomes more conservative, improving the robustness. In particular, we emphasize that our performance guarantee can match the existing results of Zhang and Li [30] by setting $c=0$. Also, when $\frac{W_{\max}}{W_{\min}}=1$, meaning that all agents’ weights are equal, our results can match those from Agrawal [1].
@@
This impossibility result generalizes known impossibility results for the unweighted case and highlights how agents’ weights fundamentally constrains achievable trade-offs. It indicates that the weighted case of facility location is more challenging than the unweighted one, as the performance guarantee of any mechanism designed for the weighted case can hardly get rid of the dependence on the input parameters $\frac{W_{\max}}{W_{\min}}$.
@@ 1.2 Related Work
The strategic FLP is a classic mechanism design problem that has been widely studied. For example, Procaccia and Tennenholtz [24] researched the approximate mechanism design without money and took this problem as an example. For single facility location in one dimension, there is a deterministic mechanism that is 2-approximate in utilitarian social cost and optimal in maximum social cost by placing the facility on the medium of all agents. This mechanism also reaches the best approximate ratio among all deterministic and strategyproof mechanisms. Zhang and Li [30] expanded this theory into weighted cases, and showed that the coordinate medium mechanism can achieve $\frac{W_{\max}}{W_{\min}}$-approximation in utilitarian social cost. Also, it is the best approximation ratio.
A lot of researches focus on this framework along with strategic FLP [9, 4]. Beyond the classic setting, extended researches include Obnoxious Facility Location (where agents seek maximal distance from facilities) [19, 12] and Heterogeneous Facility Location (where facilities can be both desirable and undesirable, from the perspective of different agents). These investigations did indeed take advantage of the category of weight, however, in their settings, the calculation of the social cost is still the sum of distances between agents and facility. In other words, agents’ weights can only be $1$ or $-1$, possessing a notable constraint. In our work, the analysis is conducted precisely by removing this constraint.
@@
Prediction-augmented algorithms, or learning-augmented algorithms in other words, is a mechanism designing method, aiming to surpass the worst-case bounds by assuming that the algorithm can predict the given instance. Lykouris and Vassilvtiskii [20] introduced consistency and robustness to analyze the performance of algorithms in this framework as the two primal metrics to measure a mechanism in this framework. After its release, many classic problems, like online paging problems [20, 18], scheduling problems [25], ski-rental problems [29, 28], and secretary problems [3] have been researched in this framework. Facility location problems are also in the list. Fotakis and Gergatsouli [15] studied online facility location problem with prediction, and Jiang and Liu [17] introduced dynamic prediction update to amend the prediction models.
However, the problems mentioned above are not strategic. In strategic problems, the predictions focus on overcoming the information limitation of privately held information rather than future information [1]. Agrawal [1] first took strategic FLP into research and designed Coordinate Medium (CM) mechanism with Prediction, also known as CMP mechanism. The CMP mechanism is simply adding $cn$ copies of $\hat{o}$ into $P$ to construct a new multiset $P^{\prime}$ to run the CM mechanism, which outputs the median of $P^{\prime}$ from each dimension separately. As is known that such median-point like mechanisms cannot be applied to general metric spaces [26], Agrawal’s work is only limited to two-dimensional Euclidean space. Our work can be regarded as the extension of their work, so we assume that agents are in two-dimensional Euclidean space, too.
@@ 2 Preliminaries
The classic FLP considers a set of agents located in a metric space (in our work, we only consider the two-dimensional Euclidean space $\mathbb{R}^{2}$). Each agent $i\in[n]$ has a preferred location $p_{i}\in\mathbb{R}^{2}$. Given the agents’ preferred location set $P=\{p_{1},\dots,p_{n}\}$, the goal of FLP is to choose a facility location $f\in\mathbb{R}^{2}$ that minimizes the average cost, measured as the average distance from all agents to the facility:
@@
where $d(\cdot,\cdot)$ denotes the Euclidean distance. This is referred to as the utilitarian cost. Given this social cost function, we can find an optimal facility location, denoted as $o(P)$.
@@
In strategic settings, agents’ real preferred locations $p_{i}$ are private information, and each agent may misreport it to reduce her individual cost $C(f,p_{i})=d(f,p_{i})$. A mechanism $\mathcal{M}$ takes reported locations $(\hat{p}_{1},\ldots,\hat{p}_{n})$ and outputs a facility location $f$. In order to prevent misreporting, we hope that the mechanism we design is strategyproof. A mechanism is said to be strategyproof if the optimal decision of any agent is reporting its real information, regardless of others’ reports. In our work, the information is the real preferred locations of agents:
@@
The most common mechanism used in strategic FLP is the Coordinate Median (CM) mechanism, which selects the facility location by independently computing the median of all coordinates (if the number of agents $n$ is even, we assume that the smaller of two medians is returned for all coordinates). It is easy to prove that this mechanism is deterministic and strategyproof in two-dimensional space, and it guarantees a bounded approximation ratio for the utilitarian objective [1]. A more general version is the Generalized Coordinate Median (GCM) mechanism. It takes the reported location set $P$ and an independent and constant point multiset $P^{\prime}$ (often called phantom points) as input, and outputs $CM(P\cup P^{\prime})$. Similarly, the GCM mechanism is also deterministic and strategyproof.
@@
We consider a prediction-augmented setting where the mechanism receives a prediction $\hat{o}\in\mathbb{R}^{2}$ of the optimal location. This prediction may comes from some of the history information of agents, and it is independent to agents’ reported locations. Mechanism designer can take advantage of the information to design a new mechanism, denoted $f(P,\hat{o})$. Also, the new mechanism needs to be strategyproof. On practice, if a strategyproof mechanism is $\gamma$-approximate, we can simply ignore the prediction, and get a prediction-augmented mechanism that is $\gamma$-consistent and $\gamma$-robust. So the challenge is to reduce the consistency guarantees, while not to sacrifice too much in robustness.
3 Prediction Augmented Mechanism and Upper
Bounds Analysis
To proof the Theorem 1, we first consider the CMP mechanism that was proved to be $\frac{\sqrt{2+2c^{2}}}{1+c}$-consistent and $\frac{\sqrt{2+2c^{2}}}{1-c}$-robust in the unweighted strategic FLP, and then try to expand it to the weighted case. We find that the CMP mechanism can match the bounds and make a great trade-off between consistency and robustness using confidence parameter $c$.

### v3

@@ Abstract
The major contribution of the current work is to provide a prediction augmented algorithmic framework for balancing the consistency and robustness over strategic agents with non-uniform weights. In particular, through a reduction technique that identifies a subset of representative instances and maps the other given locations to the representative ones, we prove that there exists a strategyproof mechanism achieving a bounded consistency guarantee of $\frac{\sqrt{(1+c)^{2}W^{2}_{\min}+(1-c)^{2}W^{2}_{\max}}}{(1+c)W_{\min}}$ and a bounded robustness guarantee of $\frac{\sqrt{(1-c)^{2}W^{2}_{\min}+(1+c)^{2}W^{2}_{\max}}}{(1-c)W_{\min}}$ in weighted settings, where $c$ can be viewed as a parameter to make a trade-off between the consistency and robustness and $W_{\min}$ and $W_{\max}$ denote the minimum and maximum agents’ weight. We also prove that there is no strategyproof deterministic mechanism that reach $1$-consistency and $O\left(n\cdot\frac{W_{\max}}{W_{\min}}\right)$-robustness in weighted FLP, even with fully predictions of all agents.
@@ 1 Introduction
Facility Location Problem (FLP) is a classic problem in operations research, computer science, and economics [8]. Its goal is to choose one or more facility locations within a given space to optimize a global cost function between users (referred to as agents) and facilities. If only one facility needs to place, we say it single FLP [^1] If not mentioned on purpose, in the following text, we say the FLP as single FLP by default.. In the problem, there are $n$ agents in a metric space $\mathcal{M}$, and each agent $i$ has her preferred location $p_{i}$ and must report a location $\hat{p}_{i}\in\mathcal{M}$ of the facility. We are required to collect this information to choose where to locate the facility $f\in\mathcal{M}$. Once we choose the facility location $f$, the individual cost of each agent $i$ is given by the distance $d(p_{i},f)$ between $p_{i}$ and $f$, where $d(\cdot)$ is the distance function defined on $\mathcal{M}$. Our objective is to optimize a global function $g(P;f)$, where $P$ is the location profile of all the agents. Naturally, here comes to a question: What kind of global cost function we should consider? Motivated by various applications in practical scenarios (see [8] for a more detailed overview), researchers in this area are often interested in the global objectives of (1) egalitarian social cost and (2) utilitarian social cost, where egalitarian social cost is to the maximum cost over all the agents, and utilitarian social cost is the average cost [1, 28]. The study of the former is more difficult as applying predictions on any existing strategyproof mechanisms is hard, for the mechanisms focus on the bounding box of $P$ that prediction can hardly get involved. So the current work is devoted to the study of the latter.
@@
Under our setting, agents are selfish, so an agent $i$ may misreport her preferred location to minimize her individual cost, which means the reported location $\hat{p}_{i}$ is not necessarily equivalent to her real preferred location $p_{i}$. This leads to the strategic facility location problem, where the focus shifts from purely optimizing a cost function to designing mechanisms that are strategyproof, i.e., mechanisms in which agents have no incentive to tell lie about their true preferences [25, 6]. This problem is first studied by Procaccia and Tennenholtz [25], focusing on the egalitarian social cost, and Dekel [11] investigates how to optimize the utilitarian social cost. Procaccia and Tennenholtz [25] prove that no deterministic and strategyproof mechanism can achieve better than a 2-approximation, when the goal is the egalitarian social cost, and Meir [22] propose the coordinate median (CM) mechanism that reaches $\sqrt{n}$-approximation in $n$-dimensional case while optimizing the utilitarian social cost. These results establish the framework for subsequent research.
Research work on FLP has been applied in establishment of public services such as schools and hospitals [16, 10]. Existing work is often based on the unweighted assumption on the agents, which means that all the agents in the environment are equivalent in importance. Although this assumption reflects equal treatment in public services, this setting is uncommon in reality, where agents often differ in importance. Therefore, the unweighted assumption rarely holds in practice. For example, in the field of healthcare, existing work suggests that the older adults are often associated with a relatively higher weights [2, 23]. Agents with distinct importance (weights) also exist in other applications including logistics planning [5], disaster assistance [7], etc. Obvliously, the unweighted setting of FLP cannot be adapted to these environments. That is why strategic FLP with non-uniform weights has attracted the attention of researchers [32, 29]. Under the weighted setting, we assume that different weight to quantifies her importance to the decision maker: agents with higher weights take greater social cost and vice versa.
@@
Existing work on strategic FLP (unweighted or weighted) usually adopts the research paradigm of worst-case analysis. This paradigm focuses on extreme scenarios that rarely occur in practice, leading to algorithm design that is overly conservative for the average case. On the other hand, average-case analysis must deal with the uncertainty in real-world problems, such as the lack of knowledge about the true distribution of agent’s preferred location in FLP. To address this issue, a novel research paradigm, prediction-augmented algorithm design has recently emerged. It allows decision-makers to access external prediction sources to obtain forecasts about the uncertainty in the problem setting, and allows them to decide how much they trust the forecasts [14, 13, 22, 32]. When the prediction is accurate, we can trust it more and get a better resulting guarantee than using worst-case analysis, i.e., consistency of the mechanism. On the other hand, when the prediction is arbitrarily wrong, we would like to ensure that the wrong prediction does not worsen the result too much and obtains a solid performance guarantee, i.e., robustness of the mechanism [20]. More precisely speaking, in the framework of prediction-augmented mechanism design of strategic FLP, given the agents’ reported locations $P$, the designer is provided a predicted location $\hat{o}$, regarding the real optimal location $o(P)$. The designer can use the information to make a mechanism, whose result is $f(P,\hat{o})$. Given a social cost function $C$, the mechanism is $\alpha$-consistent when the prediction is accurate $(\hat{o}=o(P))$ if
@@
$\max_{P,\hat{o}}{\left\{\frac{C(f(P,\hat{o}),P)}{C(o(P),P)}\right\}}\leq\beta.$
@@ 1.1 Our Contributions
We investigate the prediction-augmented mechanism design of strategic FLP in a weighted setting, where the importance of the agents is non-uniform according to the weighted values. Our goal is to design deterministic, strategyproof mechanisms that incorporate a prediction of the optimal facility location, and achieve a tunable trade-off between consistency and robustness.
@@
There exists a deterministic and strategyproof mechanism in the weighted FLP with predictions that reaches $\frac{\sqrt{(1+c)^{2}W^{2}_{\min}+(1-c)^{2}W^{2}_{\max}}}{(1+c)W_{\min}}$-consistency and $\frac{\sqrt{(1-c)^{2}W^{2}_{\min}+(1+c)^{2}W^{2}_{\max}}}{(1-c)W_{\min}}$-robustness.
$W_{\min}$ and $W_{\max}$ denote the minimum and maximum weights of agents, respectively. The mechanism uses a parameter $c\in[0,1)$ to control the degree of reliance on the prediction. When $c$ is large, the mechanism places more trust in the prediction, achieving better consistency; when $c$ is small, it becomes more conservative, improving the robustness. In particular, we emphasize that our performance guarantee matches the existing results of Zhang and Li [32] by setting $c=0$. Also, when $\frac{W_{\max}}{W_{\min}}=1$, meaning that all agents’ weights are equal, our results matches those from Agrawal [1] on the unweighted setting.
@@
Theorem 2 generalizes known impossibility results for the unweighted case and highlights how agents’ weights fundamentally constrains achievable trade-offs. It indicates that the weighted case of facility location is more challenging than the unweighted one, as the performance guarantee of any mechanism designed for the weighted case can hardly get rid of the dependence on the input parameters $\frac{W_{\max}}{W_{\min}}$.
@@ 1.2 Related Work
Strategic FLP is a classic mechanism design problem that has been widely studied. For example, Procaccia and Tennenholtz [25] research the approximate mechanism design without money and take this problem as an example. For single facility location in one dimension, there exists a deterministic mechanism that is 2-approximate in utilitarian social cost and optimal in maximum social cost by placing the facility on the medium of all agents. This mechanism also reaches the best approximate ratio among all deterministic and strategyproof mechanisms. Zhang and Li [32] expand this theory into weighted cases, and show that the coordinate medium mechanism achieves $\frac{W_{\max}}{W_{\min}}$-approximation in utilitarian social cost. Also, it is the best approximation ratio.
A lot of research work focus on this framework along with strategic FLP [9, 4]. Beyond the classic setting, extended researches include Obnoxious Facility Location (where agents seek maximum distance from facilities) [19, 12] and Heterogeneous Facility Location (where facilities can be both desirable and undesirable, from the perspective of different agents). Results in this line of work allows agents to have different weights. However, the weight of each agent is constrained in a binary set $\{-1,1\}$. In our work, the analysis is conducted precisely by removing this constraint.
@@
Prediction-augmented algorithms (also known as the learning-augmented algorithms), is an algorithm design paradigm aiming to surpass the worst-case performance guarantee by allowing the algorithm to access predictions. Lykouris and Vassilvtiskii [21] introduce consistency and robustness to analyze the performance of algorithms in this framework as the two primal metrics to measure a mechanism in this framework. After its release, many classic problems, like online paging problems [21, 18], scheduling problems [26], ski-rental problems [31, 30], and secretary problems [3] have been researched in this framework. Facility location problems are also in the list. Fotakis and Gergatsouli [15] study online facility location problem with prediction, and Jiang and Liu [17] introduce dynamic prediction update to amend the prediction models.
However, the problems mentioned above are not studied under a strategic setting. In strategic problems, the predictions focus on overcoming the information limitation of privately held information rather than future information [1]. Agrawal [1] first takes strategic FLP into research and designs Coordinate Medium (CM) mechanism with Prediction, also known as CMP mechanism. The CMP mechanism is simply adding $cn$ copies of $\hat{o}$ into $P$ to construct a new multiset $P^{\prime}$ to run the CM mechanism, which outputs the median of $P^{\prime}$ from each dimension separately. As is known that such median-point like mechanisms cannot be applied to general metric spaces [27], Agrawal’s work is conducted over the two-dimensional Euclidean space. Our work can be viewed as the extension of their work by adopting the same assumption on the metric space, while allowing agents to have non-uniform weights, which is more common in practice.
@@ 2 Preliminaries
The classic FLP considers a set of agents located in a metric space (the current work focuses on the two-dimensional Euclidean space $\mathbb{R}^{2}$). Each agent $i\in[n]$ has a preferred location $p_{i}\in\mathbb{R}^{2}$. Given the preferred location set $P=\{p_{1},\dots,p_{n}\}$, the goal of FLP is to choose a facility location $f\in\mathbb{R}^{2}$ that minimizes the average cost, measured as the average distance from all agents to the facility:
@@
where $d(\cdot,\cdot)$ denotes the Euclidean distance. This is referred to as the utilitarian cost. Given this social cost function, one can define and compute the optimal facility location, denoted as $o(P)$.
@@
In strategic settings, the real preferred locations $p_{i}$ are private information, and an agent may misreport it to reduce her individual cost $C(f,p_{i})=d(f,p_{i})$. A mechanism $\mathcal{M}$ takes reported locations $(\hat{p}_{1},\ldots,\hat{p}_{n})$ and outputs a facility location $f$. In order to prevent misreporting, we would like to ensure that the mechanism we design is strategyproof. A mechanism is said to be strategyproof if the optimal decision of any agent is reporting her real information, regardless of others’ reports. In our work, the reported information is about the real preferred locations, so the strategyproof constraints can be formulated as follows:
@@
A well known mechanism designed for strategic FLP is the Coordinate Median (CM) mechanism [1], which selects the facility location by independently computing the median of all coordinates (if the number of agents $n$ is even, we assume that the smaller of two medians is returned for all coordinates). It proves that such a deterministic mechanism is strategyproof in two-dimensional space, and it guarantees a bounded approximation ratio of $\sqrt{2}$ for the utilitarian objective [1]. The mechanism Generalized Coordinate Median (GCM) proposed by Agrawal [1] generalizes the classic version of CM by taking an independent and constant multiset $P^{\prime}$ of locations (often called the phantom points) as inputs parallel to the reported location set $P$. The location chosen by GCM for the facility is $CM(P\cup P^{\prime})$. Similarly, the GCM mechanism is also proved to be deterministic and strategyproof [1].
@@
For the weighted version of strategic FLP, we consider developing a prediction augmented mechanism that receives a prediction $\hat{o}\in\mathbb{R}^{2}$ of the optimal location. This prediction may comes from some external machine learning models trained by the history data of agents. Our mechanism can be viewed as a functions $f(P,\hat{o})$ that maps the reported preferred locations and the prediction to a location of the facility. The goal of the current paper is to ensure that the proposed mechanism is strategyproof, while achieving a competitive ratio as close to $1$ as possible when the obtained predictions are highly accurate (i.e., providing a consistency guarantee), and maintaining an acceptable competitive ratio even when the prediction error rate is high (i.e., providing a robustness guarantee).
3 Prediction Augmented Mechanism and Upper Bounds Analysis
To prove Theorem 1, we first consider the CMP mechanism that is proved to be $\frac{\sqrt{2+2c^{2}}}{1+c}$-consistent and $\frac{\sqrt{2+2c^{2}}}{1-c}$-robust in the unweighted strategic FLP, and then try to extend it to the weighted case. The major contribution here is to prove that our extended version of CMP for the weighted setting has the same upper bound on the approximation ratio with the original version when choosing a proper value for $c$, a parameter for making trade-off between the consistency and robustness.